OpenStudy (anonymous):

Calculus/physics problem? A particle is moving along the x-axis with velocity v(t)=sin2t. At time t=pi/2 it has position x = 3. Find the total distance, s traveled by this particle on the interval 0<=t<=pi.

OpenStudy (cwrw238):

if we integrate velocity we get distance integral of sin2t =- (1/2) cos2t + c x = -(1/2) cos 2t + c when x = 3 t = pi/2: so 3 = -(1/2) cos pi + c 3 = -(1/2) * -1 + c so c = 3 - 1/2 = 2.5 so relation between distance and time is x = -0.5 cos 2t + 2.5

OpenStudy (anonymous):

So would I just plug in for pi and plug in for zero and subtract the 0 value from the pi value?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

you still there? i do apologize - i gave you the wrong information in my last post

OpenStudy (cwrw238):

the x in the formula gives you the position ( or displacement) of particle on x-axis this is simple harmonic motion - the particle oscillates about a point on the axis. at time 0 the displacement x = -0.5 * cos0 + 2.5 = -0.5 + 2.5 = 2 at time pi/2 (half the period) displacement = 3 at time t displacement = -0.5 * cos 2pi + 2.5 = -0.5 * 1 + 2.5 = 2 so total distance travelled = (3-2) + (3-2) = 2 units at time pi/2 the particle is at maximum distance from center of oscillation (when cos = -1) answer distance travelled = 2

OpenStudy (anonymous):

I got that anyway

OpenStudy (cwrw238):

oh ok